OpenStudy (anonymous):

32 over 25, simplify to the square root?

OpenStudy (anonymous):

? plz be more illustrative

OpenStudy (anonymous):

Simplify the expression, square root of 32 over 25

OpenStudy (anonymous):

\[\sqrt{32/25}=\sqrt{32}/\sqrt{25} = 4\sqrt{2} / 5\]